What is the height of a triangle if the base is 60 and the hypotenuse is 61?

Answer by checkley71(8403)

assuming this is a right triangle we have:
60^2+x^2=61^2
3600+x^2=3721
x^2=3721-3600
x^2=121
x=sqrt121
x=11 answer for the third side.